OpenStudy (anonymous):

Simplify: 3a squared x 4a

OpenStudy (anonymous):

\[3a^2 * 4a = (4*3)*(a^2*a)\] \[=12a^3\]

OpenStudy (anonymous):

do u mean 3a^2 or (3a)^2?

OpenStudy (anonymous):

3a^2

OpenStudy (anonymous):

if that's the case, then mathsboy is right:)